OpenStudy (anonymous):

log10^2 +16log10^16/25+12log10^25/24+7log10^81/80+log10^25/49+2log10^7 =?.........plz note that log 10^2 is not square of 10 it means that log 2 to the base 10......

OpenStudy (anonymous):

the answer is 2 log5 + 1 plz solve for it

OpenStudy (saifoo.khan):

Can you please post a photo or something of the question?

OpenStudy (saifoo.khan):

im not sure i can understand the question

OpenStudy (saifoo.khan):

i can't see what's the base. :(

OpenStudy (anonymous):

the base for all log is 10

OpenStudy (saifoo.khan):

umm, okay.. let me see ..

OpenStudy (saifoo.khan):

http://markvestal.com/sites/default/files/images/ExpLogRules.gif Here are the rules.

OpenStudy (anonymous):

thaznnz did u solve it i solved by using these laws but i did nt got it

OpenStudy (saifoo.khan):

Umm, im trying to solve.. can you please hold on for time? i will do it in my book.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i will wait

OpenStudy (saifoo.khan):

im trying to solve but i dont get what we have to do with the numbers before log. like16, 12 etc

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (saifoo.khan):

i can't figure it out sorry.

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (anonymous):

@stormfire1 plzz help

OpenStudy (anonymous):

i think there is some confusion about the notation

OpenStudy (anonymous):

wats the doubt in question

OpenStudy (anonymous):

"log 2 to the base 10" does not make any sense is it \[\log_2(10)\] or \[\log_{10}(2)\]

OpenStudy (saifoo.khan):

second one.

OpenStudy (anonymous):

second one is the question

OpenStudy (anonymous):

\[\log(2)+16\log(16/25)+12\log(25/24)+7\log(81/80)+\log(25/49)+2\log(7)\]

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

\[\log(2)+16\log(16/25)+12\log(25/24)\] \[+7\log(81/80)+\log(25/49)+2\log(7)\]

OpenStudy (saifoo.khan):

Yes that.

OpenStudy (anonymous):

book keeping

OpenStudy (anonymous):

for example \[\log(25/49)=2\log(5)-2\log(7)\] then \(2\log(7)-2\log(7)=0\)

OpenStudy (anonymous):

i can write it all out of you like, but that is the idea. write as a bunch of different logs and combine like terms. boring exercise, but that is what you have to do

OpenStudy (anonymous):

@satellite73 7log(81/80) how to expand

OpenStudy (anonymous):

plzz one more example plz

OpenStudy (anonymous):

\[7\log(81/80)=7\log(9^2/16\times 5)=7[\log(3^4)-\log(16\times 5)]\] \[=7[4\log(3)-(\log(2^4\times 5))]\] \[=7[4\log(3)-4\log(2)-\log(5)\] \[=28\log(3)-28\log(2)-7\log(5)\] for example this sucks

OpenStudy (anonymous):

you have to do this with every term, and then add them up at the end

OpenStudy (anonymous):

\[16\log(16/25)=16(\log(2^4/5^2)=16[4\log(2)-2\log(5)]=64\log(2)-32\log(5)\] etc etc etc

OpenStudy (anonymous):

your math teacher must hate you

OpenStudy (anonymous):

why??

OpenStudy (anonymous):

to make you do all this book keeping might as well be doing your tax returns

OpenStudy (anonymous):

@satellite73 r u a student